# MultiChartQA: Benchmarking Vision-Language Models on Multi-Chart Problems

### v1

@@ 5 Conclusion
We introduce MultiChartQA, a comprehensive benchmark for evaluating multi-chart understanding using semantically related chart sets and questions spanning multiple categories. Testing 16 MLLMs reveals a significant performance gap between closed-source and open-source models. Incorporating Chain-of-Thought reasoning notably improves accuracy. MultiChartQA highlights crucial avenues for advancing MLLMs in multi-chart reasoning tasks.
@@ Acknowledgments
This research is supported by the International Summer Undergraduate Research Experience (iSURE) Program at the University of Notre Dame. We extend our gratitude to Zhuiri Xiao, Yuyang Tian, Zihao Xu, Sicheng Yang, Heng Wang, Tianyuan Yang, Yanyu Yao, Shuaicheng Ren, Bowen Zhou, Xinyang Yin, and Yuyang Bai for their participation in the human evaluation.
@@ Appendix B Error Analysis Example
Claude-3.5-Sonnet 60.08 78.20 78.48 80.13 81.35 66.41 32.22 54.23 31.58
GPT-4o 53.88 66.17 65.82 68.38 67.28 61.33 35.56 50.35 29.47
Gemini-1.5-Pro 44.97 60.15 56.96 53.85 61.93 46.88 30.00 37.32 24.21
GPT-4o-mini 36.30 50.38 59.49 37.18 42.81 43.75 13.89 38.03 18.95
@@
MiniCPM-V2.6 36.99 56.39 56.96 36.75 41.28 37.11 22.22 39.44 12.63
InternVL2-26B 38.29 57.14 54.43 47.22 54.28 43.36 10.56 40.14 15.79
InternVL-V1.5 34.20 55.64 50.63 42.74 50.92 38.67 9.44 27.82 14.74
Qwen2-VL-7B 39.53 60.15 60.76 46.79 49.24 41.41 16.67 42.96 11.58
LLaVA-OV-7B 31.94 46.62 51.90 36.54 40.83 34.38 13.33 34.15 9.47
MiniCPM-V2.5 26.96 44.36 22.78 36.75 37.46 28.52 10.00 29.93 11.58
DeepSeek-VL-7B 23.69 36.09 32.91 25.43 28.75 26.56 12.78 25.00 6.32
LLaVA-V1.6-7B 19.56 24.06 26.58 20.09 28.44 26.56 10.56 18.31 6.32
Idefics3-8B 10.75 9.02 11.39 9.40 17.13 14.06 4.44 15.85 6.32
Idefics2-8B 7.01 8.27 12.66 7.91 11.47 5.08 3.33 8.45 2.11
@@
Claude-3.5-Sonnet 68.01 78.20 79.75 73.93 78.29 69.14 63.33 60.92 45.26
GPT-4o 64.76 78.20 73.42 77.78 72.17 65.23 58.33 63.73 32.63
Gemini-1.5-Pro 53.18 63.16 62.03 61.11 71.41 52.73 51.67 41.55 25.26
GPT-4o-mini 47.95 63.91 63.29 57.26 55.50 40.62 38.89 46.13 26.32
@@
MiniCPM-V2.6 40.89 55.64 51.90 48.72 48.17 39.84 25.56 39.79 28.42
InternVL2-26B 46.05 66.17 54.43 56.20 57.65 41.02 32.22 46.13 22.11
InternVL-V1.5 38.21 54.14 48.10 45.30 50.31 31.64 26.11 38.73 18.95
Qwen2-VL-7B 35.81 55.64 59.49 39.53 47.40 36.72 12.78 36.27 13.68
LLaVA-OV-7B 26.15 33.83 39.24 29.91 33.49 29.30 10.00 30.63 12.63
MiniCPM-V2.5 25.85 36.09 29.11 29.91 30.89 27.34 12.22 26.76 22.11
DeepSeek-VL-7B 16.05 12.03 10.13 16.88 19.57 14.84 15.56 26.06 9.47
LLaVA-V1.6-7B 17.37 15.04 22.78 16.03 24.31 18.36 10.00 22.89 13.68
Idefics3-8B 12.09 18.80 15.19 15.38 22.63 7.03 3.33 13.03 7.37
Idefics2-8B 11.76 15.04 18.99 8.33 12.84 9.77 5.56 11.27 17.89
@@ Appendix G Multi-chart Examples
We demonstrate four sets of charts from four source websites.
@@
Figure 8: A chart set and the corresponding question from the OWID website.
Figure 9: A chart set and the corresponding question from the Pew website.

### v2

@@ 5 Conclusion
We introduce MultiChartQA, a comprehensive benchmark for evaluating multi-chart understanding using semantically related chart sets and questions spanning multiple categories. Testing 20 MLLMs reveals a significant performance gap between closed-source and open-source models. Incorporating Chain-of-Thought reasoning notably improves accuracy. MultiChartQA highlights crucial avenues for advancing MLLMs in multi-chart reasoning tasks.
@@ Acknowledgments
This work was supported by NSF IIS-2119531, IIS-2137396, IIS-2142827, IIS-2234058, CCF-1901059, and ONR N00014-22-1-2507. This research was supported by the International Summer Undergraduate Research Experience (iSURE) Program at the University of Notre Dame. We extend our gratitude to Zhuiri Xiao, Yuyang Tian, Zihao Xu, Sicheng Yang, Heng Wang, Tianyuan Yang, Yanyu Yao, Shuaicheng Ren, Bowen Zhou, Xinyang Yin, and Yuyang Bai for their participation in the human evaluation.
@@ Appendix B Error Analysis Example
Claude-3.5-Sonnet 65.38 76.22 74.27 72.12 77.45 71.66 36.02 70.27 42.91
GPT-4o 59.40 68.18 64.73 61.41 67.94 66.13 36.49 66.62 39.64
Gemini-1.5-Pro 53.81 64.69 61.41 55.45 62.41 55.76 32.70 58.65 35.64
GPT-4o-mini 42.89 59.09 52.70 31.21 42.48 50.00 13.27 54.86 25.82
@@
MiniCPM-V2.6 40.20 51.05 53.53 28.89 38.01 39.17 21.33 58.38 16.73
InternVL2-26B 44.38 59.79 53.11 42.63 51.63 49.77 10.43 53.78 24.73
Qwen2-VL-7B 46.99 64.34 63.90 40.61 48.87 51.15 17.06 61.08 17.09
InternVL-V1.5 42.76 58.74 56.02 41.41 49.50 49.31 10.43 51.22 17.82
LLaVA-OV-7B 36.26 51.05 43.98 30.30 35.18 38.25 13.27 52.03 13.45
MiniCPM-V2.5 30.63 40.56 26.14 31.01 33.97 32.26 10.43 45.27 15.64
DeepSeek-VL-7B 26.76 32.87 30.29 23.33 26.45 27.65 12.80 40.68 10.91
Eagle-X5-13B 26.73 34.27 38.17 20.00 25.60 25.12 10.43 37.03 13.82
LLaVA-V1.6-7B 23.06 25.52 24.48 19.80 23.83 28.80 12.80 35.14 7.64
Idefics3-8B 13.23 8.39 11.20 8.08 11.99 16.36 4.27 26.08 11.27
Idefics2-8B 9.08 4.90 9.96 6.77 8.09 8.06 3.79 17.03 9.09
Chart-Domain Specialized Models
ChartGemma 10.91 9.09 15.35 4.55 7.94 10.37 2.84 19.59 10.18
TinyChart-3B 8.82 5.24 7.88 6.36 9.72 11.06 3.79 15.68 6.91
MatCha 7.58 6.99 10.79 2.83 2.77 5.30 5.69 17.30 2.55
@@
Claude-3.5-Sonnet 69.26 73.78 71.78 68.18 72.84 71.43 65.40 73.11 56.00
GPT-4o 67.25 77.62 63.90 72.53 71.21 71.66 60.19 73.92 45.82
Gemini-1.5-Pro 59.83 68.18 61.83 63.13 67.02 59.91 54.50 62.43 41.82
GPT-4o-mini 52.44 63.99 58.51 52.53 53.69 50.46 41.71 59.59 34.18
@@
MiniCPM-V2.6 34.12 38.46 26.56 30.81 34.68 34.79 25.59 48.24 24.73
InternVL2-26B 52.18 66.43 58.92 53.54 56.24 49.54 33.18 59.86 33.45
Qwen2-VL-7B 42.78 57.34 53.94 36.26 43.33 44.47 13.74 55.54 25.09
InternVL-V1.5 45.23 58.04 47.30 38.69 47.94 44.24 26.07 57.30 30.91
LLaVA-OV-7B 31.47 43.01 31.54 25.35 29.01 31.34 10.90 50.68 15.27
MiniCPM-V2.5 25.47 31.47 16.60 20.81 24.54 24.65 12.80 43.78 16.36
DeepSeek-VL-7B 19.65 18.18 13.28 15.56 17.38 16.36 12.80 39.73 12.00
Eagle-X5-13B 28.88 33.22 31.12 25.56 26.31 29.03 12.32 46.35 15.64
LLaVA-V1.6-7B 20.37 18.18 15.77 14.85 20.57 19.82 12.32 36.35 14.91
Idefics3-8B 15.00 19.93 14.52 12.32 14.75 10.37 3.79 25.81 9.82
Idefics2-8B 13.70 13.99 14.11 6.67 8.72 11.75 7.11 24.32 13.82
Chart-Domain Specialized Models
ChartGemma 8.97 3.85 4.56 5.45 9.50 10.14 3.32 16.49 13.09
TinyChart-3B 9.75 8.74 6.22 7.07 8.65 7.14 6.64 17.43 10.55
MatCha 6.33 2.80 2.90 2.73 3.26 1.61 5.21 19.19 5.09
@@ Appendix G Multi-chart Examples
We demonstrate four sets of charts from 4 main source websites.
@@
Figure 8: A chart set and the corresponding question from the Our World in Data website.
Figure 9: A chart set and the corresponding question from the Pew Research Center website.
Appendix H Chart Source
We demonstrate the proportion distribution of 10 chart sources in Figure 10.
Figure 10: Illustrations of ten chart sources and their corresponding proportions.